OpenStudy (anonymous):

Medal A new cancer drug in clinical trials is given to a sample of 200 patients with stage 4 lung cancer over a course of 6 months. The survival rate of the patients is then tracked over 5 years. The results are as follows: After 1 year: 136 patients are still alive. After 2 years: 98 patients are still alive. After 3 years: 65 patients are still alive. After 4 years: 32 patients are still alive. After 5 years: 25 patients are still alive.

OpenStudy (anonymous):

The drug is approved by the FDA for wide-spread distribution. What is the probability that a patient with stage 4 lung cancer, who begins taking the new drug, will still be alive in three years? Express your answer as a decimal rounded to 3 decimal places. What is the probability that the patient will die in the next four years? Express your answer as a decimal rounded to 3 decimal places.

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

I try to figure out the trick, but it seems no trick here. So that P (still alive in 3 years) =65/200 = 0.325

OpenStudy (loser66):

P( still alive in 4 years) = 32/200=0.16 so that the probability that patient will die is 1 - 0.16 = 0.84

OpenStudy (anonymous):

is that for part a?

OpenStudy (loser66):

for both 2 parts. the previous post for a, next for b

OpenStudy (anonymous):

ohh thank you sooo much :)))

OpenStudy (anonymous):

is that rounded too?